I need help to lear how to solve this kind of Expression:
If 8x +5=21 then 3x +4=
How to solve it and what are those expressions called
Thank you!
Adriana

Answer by Mathtut(3670)

this is an equation and an expression....the expression is dependent upon the outcome of x in the equation.
if 8x+5=21...then we must find x .... to what 3x+4 =........
so subtract 5 from both sides
:
8x=16..divide by 2
:
x=2
:
3(2)+4=10